OpenStudy (anonymous):

1. Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2. A. y-6 = 1/2(x+3) B. y = 1/2x-6 C. y+3 = 1/2(x-6) D. x-2y = 12

OpenStudy (anonymous):

y=kx+b k is slope y=0.5x+b put point to function (6,-3) -3=0.5*6+b b=-6 answer is B

OpenStudy (anonymous):

1/2x is \(\frac{1}{2}x\) or \(\frac{1}{2x}\)? probably second because b and c is same then, so question asker should use equation editor or brackets